# Individual or collective treatments: how to target antimicrobial use to limit the spread of Mannheimia haemolytica among beef cattle?

### v1

@@ C.1 Method
In order to determine the relevant threshold for intervention, we defined a range of viable threshold values. We thus plotted both criteria on baseline scenarios (without any collective treatments). This baseline scenarios was repeated along 200 stochastic replicates in order to obtain a distribution of values for each time step. For each time step, the median value, the 5th and 95th percentiles were recorded, thus resulting in three time series for each criterion and each scenario (Figures 7,8).
@@
The interval was extracted from the HR30 Random scenario for both criteria. The upper bound was defined as the maximum point of the 95% quantile time series and the lower bound was defined by the maxium point of the median time series. The rationale behind this was to exclude values that would either be too low or too likely while also excluding any values that would be too high or too unlikely. Indeed, HR30 Random represents a baseline scenario and our selection allowed us to select only the higher half of the maximum values in this scenario. Thus, we obtained a credible range of [1,3]delimited-[]1.3[1,3][ 1,3 ] for the severity criterion, discretized with intervals of 0.2. This criterion is also characterized by parameter $\alpha$, assumed strictly lower than 0.5 to emphasize the number of individuals with severe clinical signs. This parameter was assigned a credible range of [0.1,0.4]delimited-[]01.04[0.1,0.4][ 0.1,0.4 ] with intervals of 0.1.
@@
For each threshold value in each scenario, we computed the median cumulative incidence and AMU in order to identify which threshold value had the best trade-off between case incidence and AMU.
C.2 Results
Lower thresholds were associated with higher AMU for every scenario, whatever the criterion. We chose to adopt a threshold that does not raise the AMU too greatly in a Low risk situation (HR10). Thus for the severity criterion, thresholds below 1.5 were exluded (Figure 9, top row). Among the included thresholds, 1.5 gave the lowest cumulative incidence with an $\alpha$ value of 0.4 in the HR30 and HR90 scenarios. We thus selected this threshold and parameter value (Figure 9, middle and bottom rows).
The threshold values for the slope criterion did not have a meaningful impact on the AMU in the HR10 scenarios. We thus chose to exclude values leading to a meaningful increase in AMU in the HR30 scenarios, which led to exclude values below 1 (Figure 10, middle row). We chose 1 as a threshold because it decreased the cumulative incidence in a satisfactory way while not blowing up the AMU when compared to higher threshold values.
Figure 9: Median cumulative incidence and AMU across scenarios and threshold for the severity based criterion
Figure 10: Median cumulative incidence and AMU across scenarios and threshold for the slope criterion
Threshold for the slope criterion: 1

### v2

@@ B.1 Method
In order to determine the relevant threshold for intervention, we defined a range of viable threshold values. We thus plotted both criteria on baseline scenarios (without any collective treatments). This baseline scenarios was repeated along 200 stochastic replicates in order to obtain a distribution of values for each time step. For each time step, the median value, the 5th and 95th percentiles were recorded, thus resulting in three time series for each criterion and each scenario (Figures 13,14).
@@
The interval was extracted from the HR30 Random scenario for both criteria. The upper bound was defined as the maximum point of the 95% quantile time series and the lower bound was defined by the maximum point of the median time series. The rationale behind this was to exclude values that would either be too low or too likely while also excluding any values that would be too high or too unlikely. Indeed, HR30 Random represents a baseline scenario and our selection allowed us to select only the higher half of the maximum values in this scenario. Thus, we obtained a credible range of [1,3]delimited-[]1.3[1,3][ 1,3 ] for the severity criterion, discretized with intervals of 0.2. This criterion is also characterized by parameter $\alpha$, assumed strictly lower than 0.5 to emphasize the number of individuals with severe clinical signs. This parameter was assigned a credible range of [0.1,0.4]delimited-[]01.04[0.1,0.4][ 0.1,0.4 ] with intervals of 0.1.
@@
For each threshold value in each scenario, we computed the median cumulative incidence and AMU in order to identify which threshold value had the best trade-off between case incidence and AMU. We determined the cutoff values of the cumulative incidence and AMU by computing the mean of both of the outputs accross the tested thresholds. We then selected the threshold value equal or inferior to the mean of both outputs.
B.2 Results
B.2.1 Severity criterion
In HR10 scenarios, the cumulative incidence was roughly constant across the threshold values (Figure 15). In HR30 and HR90 scenarios, it increased linearly. The incidence was superior to the cutoff in every scenario when the threshold was higher than 1.5.
Figure 15: Median and 5-95% quantiles of the cumulative incidence across scenarios and threshold values for the severity based criterion. The horizontal red line represents the mean value of the cumulative incidence accross the thresholds in the given scenario.
In HR10 scenarios, the AMU had a steady decrease for low threshold values and then plateaued to values lower than the cutoff (Figure 16). Conversely, we observed a plateau of high AMU with low threshold values in HR30 and HR90 scnearios. Thresholds values lower than 1.5 were either associated with AMU higher than the low plateau in HR10 or entering the high plateau in HR30 and HR90. Therefore, we chose 1.5 as a cutoff value for the severity criterion.
Figure 16: Median and 5-95% quantiles of the AMU across scenarios and threshold values for the severity based criterion. The horizontal red line represents the mean value of the AMU accross the thresholds in the given scenario.
B.2.2 Slope criterion
In HR10 scenarios, the cumulative incidence was also roughly constant across the threshold values (Figure 17). In HR30 and HR90 scenarios, it increased steadily. The incidence was superior to the cutoff in every scenario when the threshold was higher than 1.15.
Figure 17: Median and 5-95% quantiles of the cumulative incidence across scenarios and threshold values for the slope based criterion. The horizontal red line represents the mean value of the cumulative incidence accross the thresholds in the given scenario.
The AMU was also roughly constant in HR10 scenarios across the threshold values (Figure 18). It decreased almost linearly in HR30 and HR90 scenarios. The AMU was lower than the cutoff in every scenario when the threshold for the slope criterion was higher than 1.15. We therefore chose this value as threshold for the slope criterion.
Figure 18: Median and 5-95% quantiles of the AMU across scenarios and threshold values for the slope based criterion. The horizontal red line represents the mean value of the AMU accross the thresholds in the given scenario.
Threshold for the slope criterion: 1.15